# Reasoning Abilities of Large Language Models: In-Depth Analysis on the Abstraction and Reasoning Corpus

### v1

@@ 3.1.1. Experiment
The perceived deficiency in LLMs’ logical reasoning has been a recurrent critique, with direct attempts at solving ARC tasks yielding success rates below 10% (Mirchandani et al., 2023). To mitigate this, enhancements in LLMs’ logical reasoning are being pursued through prompting techniques like Chain of Thought (CoT) (Wei et al., 2022), Least to Most (LtM) (Zhou et al., 2023), and Tree of Thought (ToT) (Yao et al., 2023). These strategies have been shown to effectively leverage LLMs’ reasoning capabilities (Wang et al., 2019), and has the advantage of allowing for a more transparent analysis for humans, as they involve a step-by-step reasoning process. Therefore, in this experiment, we assess the impact of these prompting strategies on LLMs’ logical coherence by solving ARC tasks. We utilized advanced models, namely GPT-4 and GPT-4-32k, to test their logical reasoning with CoT, LtM, and ToT prompts.
3.1.2. Method
Simple prompts often lead LLMs to generate basic or inappropriate responses, even for straightforward question-answer tasks. To enhance LLM’s inference abilities, advanced prompting techniques such as Chain of Thought (CoT), Least to Most (LtM), and Tree of Thought (ToT) were introduced.
CoT includes examples of problem-solving processes in the prompts, encouraging LLMs to engage in causal reasoning and sequential problem-solving (Wei et al., 2022). LtM guides LLMs through a step-by-step solution process, enhancing their ability to tackle complex logic problems efficiently (Zhou et al., 2023). ToT builds on LtM by organizing solution steps into a tree structure, allowing error correction at intermediate stages by exploring different solutions (Xu et al., 2023b).
Each method has its own pros and cons. CoT lacks transparency in its process, as it presents examples all at once. LtM, while offering clearer process analysis, doesn’t evaluate steps separately, risking error accumulation. ToT provides both transparency and error reduction by considering multiple paths but is slower due to its complex evaluation process.
In our experiment, we applied all three techniques to ARC tasks, adapting them to include 2–-5 example pairs, test inputs, and general instruction for solving ARC. These instructions outline the thought process for solving representative problems. CoT’s prompts are based on human-written problem-solving instructions. LtM and ToT, however, prompt LLM to generate its own problem-solving steps. For LtM, we posed an initial question to segment the problem into manageable steps111We ask a starting question: ”If these examples follow the same rule, how would you break down the question to solve it?”, and for ToT, we implemented a similar approach but allowed for multiple proposals, incorporating additional coding to manage and evaluate these proposals in a tree structure.
3.1.3. Results
Table 2 displays the performance of LtM, CoT, and ToT when applied to 100 random tasks from the evaluation set. 222Task IDs and used prompts are available at https://bit.ly/Prompt-ARC. We conducted five iterations, reporting the percentage of correctly answered questions for each. Accuracy outside parentheses indicates correct answers only, while inside parentheses indicates both correct answers and processes, as assessed by humans.
CoT demonstrated 10.6% accuracy, while LtM and ToT exhibited around 6%. Considering both correct answers and processes, accuracy fell to 2–4%, revealing that solution processes often lacked correct reasoning, regardless of prompting technique. This suggests LLMs’ logical reasoning capabilities diverge from human reasoning.
@@ 3.1.3. Results
| 5 | 12% (6%) | 5% (2%) | 6% (2%) |
| Average | 10.6% (4.0%) | 5.6% (3.0%) | 6.4% (2.4%) |
Further analysis from three perspectives was conducted on the LLMs’ responses to ARC tasks to probe their inference mechanisms. We evaluated the tasks LLMs solved and failed, assessed cases with correct answers but flawed processes, and scrutinized the logical coherence when solving the same task type.
Solved and unsolved tasks
We conducted an analysis to identify problems that LLMs solve well and those they struggle with. Table 3 presents the accuracy of LLMs across problem difficulty levels classified by humans. The classification was based on the existing categorization relying on perceived difficulty by humans (Borsky, 2021). As a result, we discovered a tendency where problems perceived as difficult by humans align closely with those challenging for LLMs. Difficult problems shared two commonalities: 1) they required lengthy inference processes to solve, and 2) they involved considering multiple simultaneous problems to extract information about changes. An example from Fig. 10 illustrates this point: a task classified as ‘Entry’, only requires a single step of coloring, while a task classified as ‘Hard’, requires three steps: recognizing each object, identifying the priority of each object, and merging each object considering their priority. ‘Easy’ and ‘Medium’ are tasks that require relatively more complex steps than ‘Entry’ and fewer steps than ‘Hard’. Considering these observations, it can be inferred that artificial intelligence possesses simple forms of visual logic, but cannot handle complex combinations of logic.
| Entry | Easy | Medium | Hard | |
| Tasks | 2 | 20 | 46 | 14 |
| Trials | 10 | 100 | 230 | 70 |
| CoT | 100.00% | 30.00% | 0.00% | 0.00% |
| LtM | 20.00% | 19.00% | 0.00% | 2.85% |
| ToT | 50.00% | 22.00% | 0.00% | 0.00% |
| Average | 56.67% | 23.67% | 0.00% | 0.95% |
Incorrect reasoning processes
@@
Among the tasks where LLM has correctly answered, the process in some cases is flawed. Fig. 11 illustrates various cases. To solve the task, one needs to 1) identify objects within the input grid, 2) count the number of black squares inside each object, and 3) extract the object with the largest number of black squares. However, CoT, LtM, and ToT attempted to solve this task in the wrong ways. For CoT, objects appearing in the input grid were sorted first, then the object in the middle was selected. Even though CoT got the answer right, the way of sorting the object was not understandable. For LtM and ToT techniques, it was understood that a specific object needed to be selected from the given input grid to solve the task. However, they mistakenly recognized objects from the test input grid. These solutions share a commonality in that they fail to explain a logically consistent rule between the provided examples of different training inputs and outputs. In other words, regardless of prompting techniques such as CoT, LtM, or ToT, LLMs have yet to demonstrate logical consistency in discovering a single rule that applies consistently across the examples provided to solve the task.
Inconsistent task solving
If LLMs have a consistent logical structure, they should produce consistent results for the same type of task. To verify this consistency, we examined whether LLMs could solve two tasks presented in Fig. 12. Those tasks can be classified as the same type because they both require a fixed output grid size and necessitate the repetitive generation of a specific pattern. To solve these tasks, appropriate functions such as rotation and flipping are used to modify and combine the input grids. However, while all three prompt techniques could solve the first task, none could solve the second one. Considering that pattern recognition and application are relatively strong points for LLMs, the inability to consistently solve tasks at a human-like level suggests that LLMs lack a consistent logical structure.
@@ 3.1.4. Conclusion
In Section 3.1, we assessed the logical coherence of LLMs by solving 100 ARC tasks using three different prompting techniques. Our results, which showed an accuracy range from 4% to 12%, reveal variability in the reasoning performance depending on the prompting approach employed. While the LLMs displayed a rudimentary level of logical ability on simpler tasks, a deeper qualitative examination exposed underlying inconsistencies.
Our analysis of problem-solving demonstrated that LLMs could navigate certain tasks with logical precision. Yet, further scrutiny of correctly solved tasks with flawed processes unveiled gaps in their reasoning, suggesting that the LLMs’ logic could be superficial. Moreover, the observation of LLMs faltering on tasks similar to ones they had previously solved indicates a lack of robust logical structure in their reasoning.
However, it is crucial to acknowledge that this study focused on assessing logical capabilities only through varying prompting techniques. Alternative strategies such as domain-specific model fine-tuning or exploring diverse LLM architectures might yield different insights into their logical abilities and coherence. Thus, additional research and experimentation are warranted to fully understand the scope and limitations of LLMs’ inferential skills.
@@ 3.2. Capability of LLMs 2: Compositionality
In Section 3.2, we investigate compositionality, the second concept of LoTH. Compositionality refers to the ability to generate complex linguistic expressions using simpler ones. This characteristic allows individuals to effectively tackle more complex tasks by breaking sub-tasks down into simpler steps, supporting the notion that humans can solve more complex tasks when faced with them. Strong compositionality enables the resolution of complex tasks and facilitates transparent descriptions of the process, which is also an important aspect from the perspective of LLMs. This section tests compositionality by treating ARC tasks as stepwise compositions of simpler functions.
3.2.1. Experiment
@@
We test the compositional ability of LLMs through ARC. Compositionality consists of two steps: 1) understanding the meaning of the given expressions and 2) combining appropriately the expressions to obtain the desired output. Therefore, we verify whether LLMs understand the meaning of the functions provided for ARC tasks and whether they can combine the functions appropriately to produce the desired results. The result of this experiment indicates that while LLMs sufficiently understand the functions and their relationship with images, their ability to decompose and combine functions to achieve the desired outcome is weak.
3.2.2. Method
This experiment investigates the compositionality of LLMs by providing basic Domain-Specific Languages (DSLs) representing logical concepts to generate solutions for given ARC tasks. In the context of ARC, DSLs refer to functions that can solve ARC tasks. Chollet has demonstrated that humans can solve ARC using a web interface equipped with basic functions (Chollet, 2019). However, attempts to solve ARC using DSLs through computer algorithms, such as those in prior research have only achieved about 2–30% accuracy (Hodel, 2023b; Wind, 2020). This indicates that ARC tasks can be solved through combinations of DSLs, and while humans can find suitable combinations, computers seem unable to do so, which ultimately relates to the absence of compositional ability.
To analyze whether appropriate DSL is selected, we used the ToT prompting technique to generate results step by step. ToT consists of two stages: 1) generating multiple proposals for the next step and 2) selecting the most promising one. In the proposal phase, LLM receives four inputs: the current state of the ARC task, Python-formatted DSLs, natural language descriptions and examples of DSL usage, and definitions of parameters required by the DSLs. Since object information plays a significant role in ARC tasks (Xu et al., 2023a), using the Push-and-Pull clustering algorithm, additional object information based on the current state was provided. (Park et al., 2023). The form of the ARC grid is a 2D array, where object information comes in the form of (x, y) coordinates representing the position of points in the 2D array. The input DSL consists of a total of eight types: ‘rotate’, ‘flip’, ‘coloring object’, ‘rotate object’, ‘color object’, ‘move object’, ‘flip object’, ‘draw a line diagonally from a given point to the end of the grid’, ‘draw a line between two points’, and ‘color a single point’. In the evaluation phase, LLM selects four suitable DSLs to create candidates. Each of the four candidates is scored from 0.001 to 20 points, and the two candidates with the highest scores are applied to generate the next state using a separate simulator. The output is then used as an input prompt recursively provided to LLM for five iterations. The grid obtained from the final output is compared to the correct answer for evaluation.
3.2.3. Results
The experiment results are as follows: out of 99 tasks, LLM was not able to solve any tasks. The fact that humans can solve 80% of ARC tasks by combining similar DSLs implies a significant difference in compositional ability between humans and LLMs (Johnson et al., 2021). To understand the reasons for this difference, we analyzed whether 1) LLM can understand the functionality of the DSLs, and 2) LLM can find the proper combination of DSLs to solve the task.
Strong understanding of image and functions
To begin with, we analyzed whether LLM understands the functionality of the DSLs by examining its ability to accurately generate outputs that perform the specified transformations when given the DSLs and ARC grids as input. The result shows that LLM can perform the transformations on a grid without errors. This result demonstrates that LLM understands the DSLs provided as input in this experiment and knows how the state changes when applying a certain DSL.
Lack of compositionality
To determine whether LLMs can find the transformations to solve ARC tasks with combinations of DSLs, we evaluated whether the state transformations are necessary for generating the correct grid. The result indicates that while the given DSLs are insufficient to achieve the solution, they seem capable of resolving the task through subsequent steps. Furthermore, when asked to predict the next step, LLM shows a tendency to select the same DSL repeatedly, with a 75% rate of selecting the same DSL that was selected just before. This tendency indicates that LLM understands the functionality of the combined DSLs, as it can compare the resulting grid after passing through the combined DSLs with the correct grid. However, it cannot predict which DSL combinations are necessary to achieve the desired outcome.

### v2

@@ 3.1.2. Comparison Across Prompting Techniques
The perceived deficiency in LLMs’ logical reasoning has been a recurrent critique, with direct attempts at solving ARC tasks yielding success rates below 10% (Mirchandani et al., 2023). To mitigate this, enhancements in LLMs’ logical reasoning are being pursued through prompting techniques like Chain of Thought (CoT) (Wei et al., 2022), Least to Most (LtM) (Zhou et al., 2023), and Tree of Thought (ToT) (Yao et al., 2023). These strategies have been shown to leverage LLMs’ reasoning capabilities (Wang et al., 2019) effectively and have the advantage of allowing for a more transparent analysis for humans, as they involve a step-by-step reasoning process. Therefore, in this experiment, we assess the impact of these prompting strategies on LLMs’ logical coherence by solving ARC tasks.
We applied three major prompting techniques – Chain of Thought (CoT), Least to Most (LtM), and Tree of Thought (ToT) – to solve 100 ARC evaluation tasks using the GPT-4-32k model. Each technique was tested in five iterations. ARC tasks follow a few-shot learning paradigm, requiring the model to discern task rules from given example pairs and apply them to test examples. The CoT method enhances reasoning performance by utilizing a chain of thought, providing examples of ARC task solutions in the prompt. Similar contextual information was provided for LtM and ToT. LtM decomposes tasks into manageable steps and executes them sequentially, while ToT generates multiple candidates at each step post-decomposition, selecting the best candidate through voting before proceeding to the next step.
@@
| 5 | 10% (6%) | 5% (2%) | 6% (2%) |
| Average | 10.2% (4.0%) | 5.6% (3.0%) | 6.2% (2.4%) |
Comparing ARC accuracy across prompts, CoT demonstrated superior performance. Table 2 presents the results of applying LtM, CoT, and ToT to 100 randomly selected tasks from the ARC evaluation set. The experiment was repeated five times, with the percentage of correct answers included for each iteration. CoT achieved approximately 10% accuracy, while LtM and ToT showed about 6% accuracy. The superior performance of CoT may be attributed to the cumulative error propagation in ToT and LtM, where small mistakes in one step of their multi-step answer generation process can lead to compounded errors in subsequent steps. Based on these results, we exclusively used the CoT prompt in subsequent experiments.
However, when we checked the correctness of the solution process, all three prompting techniques showed low accuracy, with no significant difference at around 3%, as indicated in parentheses. These results demonstrate that while there may be differences in accuracy depending on the prompting technique, there is little variation in semantic coherence. It’s also important to note that both the results and processes fall far short of the average human accuracy of 80%. These findings suggest that LLMs lag behind humans in terms of logical coherence. To analyze the specific reasons for this, we conducted follow-up experiments. Section 3.1.3 analyzes inferential coherence, which is one aspect of logical coherence, while Section 3.1.4 examines the semantic coherence of LLMs through case studies.
3.1.3. Inferential Coherence of LLMs
In our second experiment, we designed a test to evaluate the inferential coherence of LLMs. Inferential coherence refers to the consistency in reasoning processes between the problem-solving steps and their outcomes. To assess this, we examined whether the LLM could solve problems similar to the ARC tasks it had previously solved successfully.
Fig. 5 provides an overview of the entire experiment. We began by using GPT-4o to solve test examples from 400 ARC tasks,111We used 400 tasks from the ARC training set. Unlike the previous experiment, we utilized the training set instead of the evaluation set because only the training set can be augmented through Re-ARC. repeating this process five times to identify which problems the model could consistently solve. For tasks that the LLM solved correctly at least once out of the five attempts, we employed Re-ARC (Hodel, 2024) to generate 100 additional test examples for each task. These augmented test examples are methodologically identical to the original tasks regarding their solution approach. We reasoned that if the LLM truly possesses inferential coherence and applies consistent logic in problem-solving, it should be able to solve all 100 of these augmented examples. This approach allowed us to rigorously test the LLM’s ability to generalize its problem-solving strategies across similar, but not identical, tasks.
Fig. 6 shows the results. Fig. 6(a) is a graph that represents all five iterations for the number of cumulatively correct augmented test examples. It’s important to note that the graphs for all five iterations show an exponential decrease. This graph trend doesn’t vary significantly between iterations, demonstrating low coherence regardless of the iteration. Fig. 6(b) shows the distribution of accuracy, averaged over five iterations, on 100 augmented test examples for each task. The notable points in this graph are that more than two-thirds of the tasks, 54 tasks, are concentrated below the average accuracy of 0.2, and the accuracy of the 8-th, which corresponds to the top 10%, is about 0.6. These figures demonstrate that LLMs have low inferential coherence for most ARC tasks.
3.1.4. Case Study: Semantic Coherence of LLMs
Finally, we analyzed how LLMs solved tasks in the two experiments described in Section 3.1.2 and Section 3.1.3. When evaluating not only the answers but also the process for the three prompts CoT, LtM, and ToT, we found that regardless of the prompt, the accuracy was about 3%, indicating that correct answers were being derived from incorrect processes, as shown in Fig. 7.
@@ 3.1.4. Case Study: Semantic Coherence of LLMs
To solve the task, one needs to 1) identify objects within the input grid, 2) count the number of black squares inside each object, and 3) extract the object with the largest number of black squares. However, CoT, LtM, and ToT attempted to solve this task in incorrect ways. For CoT, objects appearing in the input grid were sorted first, then the object in the middle was selected. Even though CoT got the answer right, the method of sorting the objects was not comprehensible. For LtM and ToT techniques, it was understood that a specific object needed to be selected from the given input grid to solve the task. However, they mistakenly recognized objects from the test input grid. These solutions share a commonality in that they fail to explain a logically consistent rule between the provided examples of different training inputs and outputs. In other words, regardless of prompting techniques such as CoT, LtM, or ToT, LLMs have yet to demonstrate logical coherence in discovering a single rule that applies consistently across the examples provided to solve the task.
The inconsistency of inferring correct results from incorrect processes was also observed in the second experiment conducted on the training set. Upon analyzing the natural language explanations for the 83 tasks solved at least once out of 400 training tasks, we found that in 35 of these cases, the solutions proposed by the LLM could not produce the correct answer. This finding suggests that LLMs lack semantic coherence regardless of the prompting technique or the problem at hand. In other words, LLMs are deriving outcomes unrelated to their reasoning process, as evidenced by generating correct answers from incorrect solutions.
Nevertheless, in Section 3.1.3, we identified eight tasks that the LLM could solve with an accuracy of 0.6 or higher. As shown in Fig. 8, these eight tasks consists of simple solutions such as mirroring, color mapping, and partial grid copying. These tasks shared a common characteristic of conceptual simplicity, utilizing only one of the four prior knowledge domains included in ARC: objectness, goal-directedness, numbers and counting, and basic geometry (Chollet, 2019). For the 17 tasks that required the use of two or more prior knowledge domains, the LLM failed to solve any of the 100 augmented problems. The fact that the LLM could not solve any of the augmented problems, despite having solved the original ones, suggests that LLMs are not semantically coherent and may even indicate potential data leakage.
This comprehensive analysis demonstrates that while LLMs can solve certain simple pattern recognition tasks, they struggle with more complex reasoning that requires the integration of multiple concepts. The inability to coherently apply rules across augmented test examples, coupled with the generation of correct answers through incorrect reasoning, highlights significant limitations in both the inferential and semantic coherence of current LLM systems when tackling abstract reasoning tasks like those presented in ARC.
@@ 3.1.5. Conclusion
In Section 3.1, we evaluated the logical coherence of LLMs by solving 100 ARC tasks using three different prompting techniques. Our results, showing accuracies ranging from 4% to 12%, demonstrate variability in reasoning performance depending on the prompting approach used. Additionally, when experimenting with GPT-4o on 400 training tasks, the LLM showed a high accuracy of 20%.
However, through an in-depth qualitative review, we demonstrated that the LLM’s results may not be logically coherent. For the augmented test examples (100 for each solved task), the LLM only managed to solve 8 test examples showing performance above 60%. Furthermore, we found that for 35 out of the 83 solved tasks, nearly half, of the LLM provided incorrect solution processes that could not derive the correct results. This analysis suggests that the LLM has failed to achieve human-level logical coherence.
The results of this study align with previous research asserting that logical problem-solving remains challenging for LLMs alone. One study (Wang et al., 2023a) found that LLMs can generate logically consistent reasoning with CoT prompting, even when their reasoning steps are flawed. Another study (Zhang et al., 2024a) showed that LLMs struggle with accurate self-reflection in tasks like mathematical reasoning and translation. Additionally, research (Tyen et al., 2024) revealed that LLMs often fail to detect errors in intermediate steps, exposing flaws in their reasoning process. While these studies suggest that providing more context or enforcing stronger self-reflection might improve logical reasoning (Wei et al., 2022; Wang et al., 2023a; Zhang et al., 2024a), our findings indicate that these challenges persist, suggesting the issue may not be simply a lack of information about the problem.
@@ 3.2.1. Motivation
In Section 3.2, we investigate compositionality, the second concept of LoTH.222While the Language of Thought Hypothesis principally uses the term ‘systematicity’, this study employs ‘compositionality’ as used in Fodor’s paper. This choice is made because compositionality encompasses a broader concept than systemicity. Compositionality refers to the ability to generate complex linguistic expressions given simpler ones (Fodor and Pylyshyn, 1988). This characteristic allows individuals to effectively tackle more complex tasks by breaking sub-tasks down into simpler steps, supporting the notion that humans can solve more complex tasks when faced with them. Strong compositionality enables the resolution of complex tasks and facilitates transparent descriptions of the process, which is also an important aspect from the perspective of LLMs.
@@
This section uses ARC to test the compositionality of LLMs. There are previous studies that have tested a model’s compositionality by providing functions in the prompt that can be combined to solve tasks, and then checking if the model can solve them (Sinha et al., 2024). Similarly, in this study, we also provide step-by-step functions, which we refer to as DSL (Domain Specific Language), and then conduct experiments to verify whether they can solve ARC tasks. Additionally, to understand why tasks might not be solved, we conducted further experiments on the model’s comprehension of these functions. Therefore, we verify whether LLMs understand the meaning of the functions provided for ARC tasks and whether they can combine the functions appropriately to produce the desired results. The result of this experiment indicates that while LLMs sufficiently understand the functions and their relationship with images, their ability to decompose and combine functions to achieve the desired outcome is weak.
3.2.2. Compositionality of LLMs
In the first experiment, to measure compositionality, we provided LLM with information about DSL and asked them to solve given ARC tasks. Fig. 9 illustrates the structure of the entire experiment. If an LLM possesses sufficient compositionality, it should be able to select appropriate DSLs and their arguments for a given goal. However, in cases where the LLM failed to choose the correct DSL, we divided the conditions further to identify the cause. These conditions were whether the LLM understood the goal and whether it understood the solution process. To analyze the results according to each condition, four types of experiments were conducted: 1) given only DSL, 2) given correct output along with DSL, 3) given human descriptions to ARC test examples along with DSL, and 4) given both correct output grid and human descriptions along with DSL. Providing the correct output grid demonstrates compositionality based on knowing or not knowing the goal while providing human descriptions shows the impact of natural language descriptions on compositionality. All experiments were repeated 10 times each for 158 tasks.333The 158 tasks correspond to those that can be solved within 10 steps using the given DSL, out of the total 800 publicly available ARC tasks. Lastly, to establish a baseline, we also conducted human experiments. Seven participants were constrained to solve the tasks using only the same DSLs given to the LLMs.
Each DSL was given as a Python function. In this experiment, we used 19 types of DSL capable of solving ARC tasks. The prompts commonly included a brief explanation of ARC, DSL function code with comments, DSL usage examples, demonstration examples of tasks, inputs for the test examples, and object information of the test inputs. Object information is one of the crucial parameters in solving ARC tasks, which is why it was added to the prompt. We used the PnP algorithm (Park et al., 2023) to extract object information from ARC tasks. The LLM returned a JSON-formatted string representing the chosen DSL and arguments at each step, which was used to verify whether the LLM reached the correct test output with an appropriate combination of DSL and arguments. We used the most recent model, GPT-4o, for this experiment.
| w/o Human Description | w/ Human Description | Accuracy of Human | |
| w/o Test Output | 3% (0.93) | 8% (0.97) | 86% |
| w/ Test Output | 9% (0.96) | 14% (0.96) | X |
The experimental results are shown in Table 3. An average accuracy of 9% was observed when the test output was provided, and an average accuracy of 3% was observed without the test output. Similarly, it is an interesting finding that compositionality strengthens when human explanations are included in the prompt. As shown in Table 3, this improvement occurs at a similar rate to the test output. Cronbach’s alpha measurements for each experiment showed consistency in responses, with all four experiments scoring above 0.7. However, the series of results fell significantly short of the 86% accuracy corresponding to human performance. These low accuracy rates suggest issues in the LLM’s DSL composition process.
3.2.3. Analysis of compositional failures resulting from DSL misinterpretation
The issue is that the average accuracy described in Table 3 doesn’t solely reflect compositionality. When we use a DSL to solve ARC tasks, we can think about the likelihood of choosing the right DSL for each step in two parts: 1) How well LLMs understand the DSL: This is reflected in how accurately it can predict the next grid when given the DSL instructions. 2) How necessary each predicted grid is in creating the final solution: This relates to how well the steps fit together to solve the task. The overall chance of picking the correct DSL for all steps depends on both of these factors working together. To solve a task, all DSLs must be correct for 10 steps. Reflecting this, we can estimate as shown in Eq. 1 below. In this equation, represents the number of steps, represents the number of tasks at step , represents the single-step accuracy, and represents the difficulty of composition for each task. We assumed that the LLM’s compositionality could vary depending on the information provided to the LLM and the task.
To determine the task accuracy considering only the compositional difficulty, we must estimate the value when . Therefore, we conducted an additional experiment, as shown in Fig. 10 to verify the probability of not finding an appropriate DSL due to the inability to predict the output grid when selecting a DSL.
In the additional experiment, we checked how accurately the LLM could generate the correct output grid when given the DSL and ARC input grid. The experiment was conducted on 158 tasks, with each task repeated 10 times. The correct DSL and its argument chain for ARC tasks, created by humans while solving the tasks, were provided to the LLM. We prioritized using the solution with the shortest step length among the solutions provided by the human solvers. If the LLM can predict all subsequent output grids when given a specific DSL, it should be able to correctly produce the output grid regardless of the step length, since both the input grid and DSL were provided.
The LLM’s accuracy according to the number of DSL steps needed to solve the task is shown in Fig. 11. We can observe a tendency for accuracy to decrease as the number of steps required to solve the task increases. From the above results, we used a weighted average to calculate in Eq. 2. Eq. 2 shows the formula used to calculate p. In this equation, represents the single-step accuracy, represents the number of tasks at step , and represents the accuracy at step . Based on this, we estimated the single-step accuracy to be 81%.
Below Table 4 shows the expected modified accuracy when p corresponding to 0.8 is adjusted to 1. This table reflects only compositionality, confirming that nearly 30% of tasks can be solved when given the test output and human description. It should be noted that when both the correct answer and the natural language description leading to the answer are provided, there is a consistent increase of about 10% point. This suggests that each element positively influences the average compositional difficulty of the problem, represented as in Eq. 1.
| w/o Human Description | w/ Human Description | |
| w/o Test Output | 5% | 15% |
| w/ Test Output | 17% | 29% |
3.2.4. Case Study: Enhancement of Compositionality through Human Descriptions
One notable observation was the enhanced compositionality when human descriptions of problem-solving methods were included in prompts. To investigate how LLMs could solve tasks with human descriptions, we analyzed the solution processes of 13 additional tasks solved when human descriptions were provided. Results indicate that human descriptions facilitate task input abstraction and action abstraction, thereby improving problem-solving capabilities. For instance, as illustrated in Fig. 12, LLMs fail to recognize patterns in the correct output without descriptions; however, they immediately identify patterns such as an ‘X’ shape with descriptions. These findings suggest the potential to enhance LLMs’ reasoning performance by incorporating abstracted task information.